# Lu Gan, Xi Li Independent Researchers

### v1

@@ Abstract
The development of high-performance, on-device keyword spotting (KWS) systems for ultra-low-power hardware is critically constrained by the scarcity of specialized, multi-command training datasets. Traditional data collection through human recording is costly, slow, and lacks scalability. This paper introduces SYNTTS-COMMANDS, a novel, multilingual voice command dataset entirely generated using state-of-the-art Text-to-Speech (TTS) synthesis. By leveraging the CosyVoice 2 model and speaker embeddings from public corpora, we created a scalable collection of English and Chinese commands. Extensive benchmarking across a range of efficient acoustic models demonstrates that our synthetic dataset enables exceptional accuracy, achieving up to 99.5% on English and 98% on Chinese command recognition. These results robustly validate that synthetic speech can effectively replace human-recorded audio for training KWS classifiers. Our work directly addresses the data bottleneck in TinyML, providing a practical, scalable foundation for building private, low-latency, and energy-efficient voice interfaces on resource-constrained edge devices.
@@ 1 Introduction
Our core contribution is the demonstration that synthetic data is not merely a substitute but a superior alternative for KWS model training. We validate SYNTTS-COMMANDS by training a spectrum of efficient acoustic models. The results are striking: top-performing models achieve over 99.4% accuracy on English and nearly 98% on Chinese commands. This performance firmly establishes that TTS-generated datasets can effectively replace traditional, labor-intensive speech collections. By providing this scalable, high-quality data foundation and validating its efficacy, we aim to accelerate the development of more private, efficient, and intelligent on-device voice interfaces, thereby pushing the frontiers of TinyML forward.
@@ 3.1 Source Data Collection
3.2 Command Line Design
@@
The English command set includes:
@@
The Chinese command set provides equivalent functionality:
@@ 3.3 Audio Selection and Quality Filtering
To ensure high-quality voice command samples for our CosyVoice-Commands dataset, we implemented a rigorous selection process from both the VoxCeleb1, VoxCeleb2, and Free ST Chinese Mandarin Corpus datasets. For each speaker, we analyzed all available audio files across multiple quality metrics including signal-to-noise ratio (SNR), speech ratio, duration, and a composite quality score. Table 1 show the comparative statistics of audio features across the datasets used in our study.
@@ 5.1 Benchmark Model Selection and Rationale
To validate the effectiveness and utility of the proposed SYNTTS-COMMANDS Dataset, we conducted a comprehensive benchmark evaluation using a suite of state-of-the-art acoustic models. The primary objectives of this evaluation are twofold: (1) to empirically demonstrate that the dataset supports high-accuracy classification of voice commands in both English and Chinese, thereby confirming its quality and effectiveness; and (2) to establish a foundational performance baseline for future research in voice-command recognition for resource-constrained smart devices.
@@ 5.2 Benchmark Results and Analysis
We present a comprehensive benchmark of six representative acoustic models on the SYNTTS-COMMANDS Dataset across both English (EN) and Chinese (ZH) subsets, with results summarized in Table 7. All models are evaluated in terms of classification accuracy, cross-entropy loss, and parameter count, providing insights into the trade-offs between performance and model complexity in multilingual voice command recognition.
Table 7: Benchmark Results on the SYNTTS-COMMANDS Dataset (EN and ZH)
@@
Our results demonstrate that the SYNTTS-COMMANDS dataset supports high-accuracy command recognition in both languages. Notably, the top-performing models achieve over 99.4% accuracy on English and nearly 98% on Chinese, confirming the dataset’s quality and suitability for real-world deployment. Among all models, CRNN attains the best English accuracy (99.50%) and the lowest loss (0.0163), while MobileNet-V1 yields the lowest loss on Chinese (0.0552) and competitive English performance (matching CRNN’s 99.50% accuracy). Interestingly, EfficientNet shows slightly higher Chinese accuracy (97.93%) than MobileNet-V1 (97.92%), despite a higher loss—suggesting better calibration or robustness in its predictions.
In contrast, lightweight models exhibit a clear accuracy–complexity trade-off. MicroCNN, with only  4.2K parameters, achieves 93.22% accuracy on English but drops significantly on Chinese (80.14%), highlighting the increased difficulty of modeling tonal and phonetic richness in Mandarin with ultra-compact architectures. DS-CNN and TC-ResNet, with under 70K parameters, already recover strong performance (>>96.5% in both languages), underscoring their efficiency for resource-constrained applications.
An intriguing observation is the parameter asymmetry between language-specific deployments: MobileNet-V1 uses 4.31M parameters in English but only 2.65M in Chinese. This stems from differences in the number of output classes (e.g., distinct command vocabularies or phoneme sets) between the two subsets, illustrating how model size can adapt to linguistic complexity.
@@ 6 Discussion
Our benchmark results on the SYNTTS-COMMANDS dataset validate a significant shift in speech data creation: synthetic speech generated by modern TTS systems has matured to the point where it can effectively replace—and in some aspects surpass—human-recorded data for training voice command classifiers. Unlike traditional data collection, which involves lengthy processes of speaker recruitment, audio recording, and manual annotation, our approach enables rapid, scalable generation of linguistically diverse and acoustically consistent training samples. This methodology not only drastically reduces cost and time, but also offers fine-grained control over speaker characteristics and acoustic conditions—effectively addressing a key bottleneck in TinyML development.
@@
It should be noted that the current evaluation focuses exclusively on in-class command recognition, where all test samples belong to a predefined command set. This design intentionally isolates the core research question—whether synthetic commands can reliably replace real speech during model training. For real-world use, however, a voice interface must also handle out-of-vocabulary inputs and environmental noise. While the current version of SYNTTS-COMMANDS does not include negative samples, developers can readily augment it with public audio datasets such as ESC-50 [piczak2015dataset] and UrbanSound8K [salamon2014dataset] to improve rejection capabilities. Such a hybrid approach maintains the scalability of synthetic commands while incorporating real-world acoustic variability.

### v2

@@ Abstract
The development of high-performance, on-device keyword spotting (KWS) systems for ultra-low-power hardware is critically constrained by the scarcity of specialized, multi-command training datasets. Traditional data collection through human recording is costly, slow, and lacks scalability. This paper introduces SYNTTS-COMMANDS, a novel, multilingual voice command dataset entirely generated using state-of-the-art Text-to-Speech (TTS) synthesis. By leveraging the CosyVoice 2 model and speaker embeddings from public corpora, we created a scalable collection of English and Chinese commands. Extensive benchmarking across a range of efficient acoustic models demonstrates that our synthetic dataset enables exceptional accuracy, achieving up to 99.5% on English and 98% on Chinese command recognition. These results robustly validate that synthetic speech can effectively replace human-recorded audio for training KWS classifiers. Our work directly addresses the data bottleneck in TinyML, providing a practical, scalable foundation for building private, low-latency, and energy-efficient voice interfaces on resource-constrained edge devices. The dataset and source code are publicly available at https://github.com/lugan113/SynTTS-Commands-Official.
@@ 1 Introduction
Our core contribution is the demonstration that synthetic data is not merely a substitute but a superior alternative for KWS model training. We validate SYNTTS-COMMANDS-MEDIA by training a spectrum of efficient acoustic models. The results are striking: top-performing models achieve over 99.4% accuracy on English and nearly 98% on Chinese commands. This performance firmly establishes that TTS-generated datasets can effectively replace traditional, labor-intensive speech collections. By providing this scalable, high-quality data foundation and validating its efficacy, we aim to accelerate the development of more private, efficient, and intelligent on-device voice interfaces, thereby pushing the frontiers of TinyML forward.
@@ 3.1 Source Data Collection
3.2 Media Command Line Design
@@
The English media command set includes:
@@
The Chinese media command set provides equivalent functionality:
@@ 3.3 Audio Selection and Quality Filtering
To ensure high-quality voice command samples for our Media dataset, we implemented a rigorous selection process from both the VoxCeleb1, VoxCeleb2, and Free ST Chinese Mandarin Corpus datasets. For each speaker, we analyzed all available audio files across multiple quality metrics including signal-to-noise ratio (SNR), speech ratio, duration, and a composite quality score. Table 1 show the comparative statistics of audio features across the datasets used in our study.
@@ 5.1 Benchmark Model Selection and Rationale
To validate the effectiveness and utility of the proposed SYNTTS-COMMANDS-MEDIA Dataset, we conducted a comprehensive benchmark evaluation using a suite of state-of-the-art acoustic models. The primary objectives of this evaluation are twofold: (1) to empirically demonstrate that the dataset supports high-accuracy classification of voice commands in both English and Chinese, thereby confirming its quality and effectiveness; and (2) to establish a foundational performance baseline for future research in voice-command recognition for resource-constrained smart devices.
@@ 5.2 Benchmark Results and Analysis
We present a comprehensive benchmark of six representative acoustic models on the SYNTTS-COMMANDS-MEDIA Dataset across both English (EN) and Chinese (ZH) subsets, with results summarized in Table 7. All models are evaluated in terms of classification accuracy, cross-entropy loss, and parameter count, providing insights into the trade-offs between performance and model complexity in multilingual voice command recognition.
Table 7: Benchmark Results on the SYNTTS-COMMANDS-MEDIA Dataset (EN and ZH)
@@
Our results demonstrate that the SYNTTS-COMMANDS-MEDIA dataset supports high-accuracy command recognition in both languages. Notably, the top-performing models achieve over 99.4% accuracy on English and nearly 98% on Chinese, confirming the dataset’s quality and suitability for real-world deployment. Among all models, CRNN attains the best English accuracy (99.50%) and the lowest loss (0.0163), while MobileNet-V1 yields the lowest loss on Chinese (0.0552) and competitive English performance (matching CRNN’s 99.50% accuracy). Interestingly, EfficientNet shows slightly higher Chinese accuracy (97.93%) than MobileNet-V1 (97.92%), despite a higher loss—suggesting better calibration or robustness in its predictions.
In contrast, lightweight models exhibit a clear accuracy–complexity trade-off. MicroCNN, with only $\approx$4.2K parameters, achieves 93.22% accuracy on English but drops significantly on Chinese (80.14%), highlighting the increased difficulty of modeling tonal and phonetic richness in Mandarin with ultra-compact architectures. DS-CNN and TC-ResNet, with under 70K parameters, already recover strong performance (>>96.5% in both languages), underscoring their efficiency for resource-constrained applications.
@@ 6 Discussion
Our benchmark results on the SYNTTS-COMMANDS-MEDIA dataset validate a significant shift in speech data creation: synthetic speech generated by modern TTS systems has matured to the point where it can effectively replace—and in some aspects surpass—human-recorded data for training voice command classifiers. Unlike traditional data collection, which involves lengthy processes of speaker recruitment, audio recording, and manual annotation, our approach enables rapid, scalable generation of linguistically diverse and acoustically consistent training samples. This methodology not only drastically reduces cost and time, but also offers fine-grained control over speaker characteristics and acoustic conditions—effectively addressing a key bottleneck in TinyML development.
@@
It should be noted that the current evaluation focuses exclusively on in-class command recognition, where all test samples belong to a predefined command set. This design intentionally isolates the core research question—whether synthetic commands can reliably replace real speech during model training. For real-world use, however, a voice interface must also handle out-of-vocabulary inputs and environmental noise. While the current version of SYNTTS-COMMANDS-MEDIA does not include negative samples, developers can readily augment it with public audio datasets such as ESC-50 [piczak2015dataset] and UrbanSound8K [salamon2014dataset] to improve rejection capabilities. Such a hybrid approach maintains the scalability of synthetic commands while incorporating real-world acoustic variability.